# “It Explains What I am Currently Going Through Perfectly to a Tee”: Understanding User Perceptions on LLM-Enhanced Narrative Interventions

### v1

@@ 5.1.2. RQ2: LLM-Enhanced Stories as an Intervention to Help People Manage Negative Thoughts
Our findings suggest that LLM-enhanced stories were relatively more effective in helping individuals reduce their belief in their negative thoughts, aligning with prior literature that underscores the benefits of structured reflection activities for mental health management (O’Leary et al., 2018; Bhattacharjee et al., 2024a, 2022). These stories facilitated reflection on personal thought patterns, allowing participants to draw parallels between their own challenges and those faced by the characters within the narratives (Akram et al., 2020). The LLM-enhanced stories appeared to promote diffuse sociality (Burgess et al., 2019), reducing feelings of isolation by showing that others may also experience similar struggles. Participants noted that this sense of connection reassured them that they were not alone in their challenges.
Narratives that depicted progress, even if minor, provided participants with hope and a sense of direction. Many participants utilized the stories as an opportunity to critically assess and reevaluate their thought patterns. This reflection encouraged them to consider possible changes such as adopting new perspectives on their experiences and exploring alternative actions (Bhattacharjee et al., 2024a), aligning with previous research on the benefits of peer narratives (Bhattacharjee et al., 2022; Akram et al., 2020; Lipsey et al., 2020).
@@
The LLM-enhanced stories also highlight the role of affective empathy in DMH interventions (Sharma et al., 2023a, 2021; Picard, 2000). Affective empathy describes an entity’s ability to perceive, respond to, and convey emotions (Picard, 2000). The LLM-enhanced stories demonstrated this capability by dynamically adapting the content to reflect how users felt about the challenges they encountered. Participants particularly valued how the stories acknowledged the characters’ struggles, which resonated deeply with their own experiences. They reported significant benefits, such as the ability to reevaluate feelings of failure or explore alternative solutions — outcomes that align with the recognized advantages of empathetic interactions in DMH tools (Sharma et al., 2023b; Ardenghi et al., 2024; Pereira and Díaz, 2019).
@@
Finally, relatability was often cited as a key factor enhancing the impact of the LLM-enhanced stories within our study, underscoring the importance of tailoring DMH content to individual contexts (Bhattacharjee et al., 2023b; Baumel et al., 2019). Participants frequently noted that the stories reflected their own experiences, aiding them in visualizing how the advice could be applicable to their own lives (Bhattacharjee et al., 2022). However, excessive relatability detracted from the story’s applicability, as it gave the impression of the story being unrealistically tuned to an individual’s unique situations. These comments bear similarities with the uncanny valley effect (Seyama and Nagayama, 2007) encountered in robotics, which entails discomfort when replicas appear almost, but not quite, human. In the context of DMH, excessively tailored stories can also seem similarly artificial, potentially causing participants to view them as forced and potentially undermining their effectiveness.
@@ 5.2.2. Refining Tone and Wording of AI-generated Content
While LLM-enhanced stories demonstrated promising potential as narrative interventions, some participants expressed concerns about the writing style and tone, which they felt diminished the overall impact. This issue arose from the use of certain phrasing or the participants’ familiarity with the way ChatGPT responds, making the stories feel less authentic. The tone of the responses was frequently mentioned, with some participants finding the tone to be too mechanical lacking the personal touch necessary for conversations around managing negative thoughts (Cheng et al., 2022). Others noted that the overly optimistic tone in certain stories did not always align with or acknowledge the severity of their struggles, which, in turn, made them feel that their challenges were being minimized (Haugeland et al., 2022).
@@ 5.2.3. Structured Reflection Activities
Participants in our study expressed appreciation for the reflection activity embedded within the stories, noting that such exercises helped them reevaluate their thoughts and emotions and potentially identify solutions to their problems. However, the reflection activity in our study was limited to a single question. Future studies might consider drawing inspiration from guided reflection activities, which involve a sequence of targeted questions focusing on specific aspects of one’s thoughts (O’Leary et al., 2018). These multi-question formats allow for a more structured presentation of thoughts, capturing nuances that a single question might miss. Incorporating such detailed questioning can enhance the benefits of LLM-enhanced stories by providing deeper insights into their thoughts and challenges, and potentially helping find ways of solving them. (Bhattacharjee et al., 2024a; Sharma et al., 2023b). However, it is important to note that answering these more detailed questions requires more time and effort from users. Therefore, careful consideration should be given to the timing and frequency of these questions to maintain a balance between depth of engagement and participant burden.
@@ 5.2.4. Improving Peer Support and Crowd-Powered Platforms
In the past, peer support and crowd-powered platforms have shown promising potential in providing wellbeing support, yet they face several inherent challenges that could be alleviated by an approach like ours. For example, the act of sharing and writing content can be daunting; LLMs can assist crowdworkers and peers by providing example stories and offering writing support to edit and restructure users’ own experiences (Kim et al., 2024; Reza et al., 2023). Additionally, some individuals may be reluctant to open up to real people. In these cases, the ability of LLMs to craft highly personalized narratives in a completely private manner — without the need to share personal details — could significantly increase their appeal, particularly for more inhibited users. Nonetheless, it is crucial to position LLMs as supportive tools that enhance the human-driven process and maintain the essential human element in peer support intact.
@@ 5.2.5. Considerations for People’s Attitude Toward AI
While our study did not specifically measure attitudes towards AI, some participants’ comments suggest that such attitudes may influence their perceptions of LLM-enhanced stories. Some of our participants expressed resistance to the idea of AI-based tools providing support for managing negative thoughts while others did not, suggesting a wide range of expectations regarding AI’s role in personal interactions (kaya2024roles; Liu et al., 2022). This variation suggests that future research could benefit from directly measuring attitudes toward AI (Grassini, 2023) and examining how these attitudes shape preferences for LLM-generated stories. Additionally, future DMH tools could incorporate a feature allowing users to express their attitudes toward AI during registration. This mechanism could enable the system to adjust its responses to better match individual expectations and comfort levels with AI technology. Moreover, AI systems could dedicate time to build trust gradually, rather than immediately offering wellbeing support, which some individuals might find uncomfortable (O’Dell et al., 2022).
5.3. Limitations and Future Works
Our study has several limitations that merit attention. First, the engagement of participants with the stories was limited to a one-time survey activity. While brief reflection activities like these are often tested in singular instances, the long-term effects of such interventions are also critical (Bhattacharjee et al., 2024a). A longitudinal deployment could uncover factors such as time, user mood, emotion, and social context that might influence engagement with the stories (Bhattacharjee et al., 2023b). Future research should explore these dimensions.
@@ 5.3. Limitations and Future Works
Secondly, our study population was confined to young adults. While some findings from this demographic might be generalizable to other groups, conducting similar studies with different populations could reveal new insights into their preferences and responses. Future studies should also consider including individuals with clinically elevated levels of depression and anxiety symptoms, keeping in mind the heightened need for risk management processes and oversight in such research.
@@
Lastly, our study relied primarily on user perceptions of the LLM-enhanced stories — their views on various qualities of the stories and their effectiveness as interventions. That is, we did not evaluate objective qualities of the content itself. While we constrained the LLMs to work within the framework of previously established narratives supporting evidence-based psychological strategies, future research should assess adherence to these strategies and other inappropriate content that may appear in generated stories, potentially involving clinicians in the evaluation process to ensure the accuracy and appropriateness of the content.
@@ 6. Conclusion
LLMs offer promising capabilities to dynamically adapt and generate stories that can resonate with the unique challenges individuals face. Motivated by this potential, our study leveraged LLM to enhance existing human-written stories from past DMH interventions, comparing the LLM-enhanced versions with their original counterparts. We conducted an experiment with young adults, including crowdworkers and students, and observed that LLM-enhanced stories were perceived to do better than human-written ones in conveying key takeaways, promoting reflection, and reducing belief in negative thoughts. These narratives were not only perceived as more relatable but also maintained a level of authenticity comparable to human-written stories. Our findings illuminate potential areas for future work on LLM-enhanced narrative interventions, such as balancing relatability with avoiding implausibility and refining the AI’s tone and writing style. We believe our work marks a significant step towards enhancing DMH interventions with emerging AI technology, enabling them to dynamically adapt based on user challenges.
@@ Acknowledgements.
This work was supported by grants from the National Institute of Mental Health (K01MH125172, R34MH124960), the Office of Naval Research (N00014-18-1-2755, N00014-21-1-2576), the Natural Sciences and Engineering Research Council of Canada (RGPIN-2019-06968), and the National Science Foundation (2209819). In addition, we acknowledge a gift from the Microsoft AI for Accessibility program to the Center for Behavioral Intervention Technologies that, in part, supported this work (http://aka.ms/ai4a). Ananya Bhattacharjee would also like to acknowledge support from the Inlight Research Fellowship, the Schwartz Reisman Institute Graduate Fellowship, and the Wolfond Scholarship in Wireless Information Technology.
Appendix A Examples of Stories
@@
Please answer this reflective question: Can you think of a time when reaching out to someone helped you break out of a rut? Relying on others may sometimes make you feel vulnerable, but sometimes it is in this vulnerability that we find our greatest strengths. There’s no shame in seeking help, after all, it’s okay to not have all the pieces together. We’re all a work in progress.
@@
I was thrust into independence at an early age, having realized that relying on others often led to disappointment. This developed a hardened shell around me, especially in romantic relationships where the fear of vulnerability reigned. I yearned for connection, yet the thought of opening up was paralyzing. Once I did open up, the outcome was a heartbreak that took me seven years to recover from.
@@
During those years, a shroud of chronic depression enveloped me. I pushed everyone away, building an even thicker protective barrier around myself.
@@
I feel too shy or anxious when I’m in social situations I feel overwhelmed and awkward in social situations, like I don’t know how to fit in or what to say. The anxiety just takes over, and I end up feeling invisible or out of place. I am a shy person. It’s part of who I am. I sometimes feel bad about this, like I have lost out on experiences because of being shy and not showing my true self to the world. I’ve often found myself tangled in the webs of social anxiety, like a silent character adrift in a sea of voices. There’s this persistent whisper in my psyche, convincing me that everyone in the room is scrutinizing each slip, each stumble, and each stutter.
It’s exhausting, and I wish I could relax and be myself around others. While I know I am as worthy of having a voice as anyone else, there’s this worry in my head that others will judge me.

### v4

@@ 5.1.2. RQ2: Personalized LLM-Enhanced Stories as an Intervention to Help People Manage Negative Thoughts
Our findings suggest that LLM stories were relatively more effective in helping individuals reduce their belief in their negative thoughts, aligning with prior literature that underscores the benefits of structured reflection activities for mental health management (O’Leary et al., 2018; Bhattacharjee et al., 2024a, 2022). These stories facilitated reflection on personal thought patterns, allowing participants to draw parallels between their own challenges and those faced by the characters within the narratives (Akram et al., 2020). The LLM stories appeared to promote diffuse sociality (Burgess et al., 2019), reducing feelings of isolation by showing that others may also experience similar struggles. Participants noted that this sense of connection reassured them that they were not alone in their challenges.
@@
The LLM stories also highlight the role of affective empathy in DMH interventions (Sharma et al., 2023a, 2021; Picard, 2000). Affective empathy describes an entity’s ability to perceive, respond to, and convey emotions (Picard, 2000). The LLM stories demonstrated this capability by dynamically adapting the content to reflect how users felt about the challenges they encountered (Motahar et al., 2024). Participants particularly valued how the stories acknowledged the characters’ struggles, which resonated deeply with their own experiences. They reported significant benefits, such as the ability to reevaluate feelings of failure or explore alternative solutions — outcomes that align with the recognized advantages of empathetic interactions in DMH tools (Sharma et al., 2023b; Ardenghi et al., 2024; Pereira and Díaz, 2019).
@@
We also noticed that relatability was often cited as a key factor enhancing the impact of the LLM stories within our study, underscoring the importance of tailoring DMH content to individual contexts (Bhattacharjee et al., 2023b; Baumel et al., 2019; Gatos et al., 2021). Participants frequently noted that the stories reflected their own experiences, aiding them in visualizing how the advice could be applicable to their own lives (Bhattacharjee et al., 2022). However, excessive relatability detracted from the story’s applicability, as it gave the impression of the story being unrealistically tuned to an individual’s unique situations. These comments bear similarities with the uncanny valley effect (Seyama and Nagayama, 2007) encountered in robotics, which entails discomfort when replicas appear almost, but not quite, human. In the context of DMH, excessively tailored stories can also seem similarly artificial, potentially causing participants to view them as forced and potentially undermining their effectiveness.
Finally, we emphasize that LLM-based narrative interventions are not necessarily intended to be implemented in the exact form used in this study, nor are they alone sufficient to address broader wellbeing challenges. Narrative-based approaches are increasingly integrated into interventions targeting specific goals, such as managing depression and anxiety (Rodríguez Vega et al., 2014; Meyerhoff et al., 2024), promoting physical activity (Saksono et al., 2021), or encouraging healthy habits (Grimes et al., 2008). Interventions like ours can function as modular components within a broader framework, addressing specific needs while complementing other strategies. For instance, the current version of our SSI aims to communicate psychological principles for managing commonly experienced negative thoughts, such as feeling down on oneself, worrying about the future, or feeling let down by others. Similar methods could be used to reframe everyday stressors (Bhattacharjee et al., 2024a; Meyerhoff et al., 2024) in a way that offers users brief, supportive narratives as a way to regulate immediate emotional responses. These SSIs could be integrated into larger DMH programs, where they can coexist with other approaches such as skill-building exercises or psychoeducational modules. However, deploying LLM-generated narratives for severe conditions, such as post-traumatic stress disorder (PTSD) or suicidal ideation, would likely require additional safeguards such as direct clinician involvement and supervision prior to story delivery.
@@ 5.2.2. Refining Tone and Wording of AI-generated Content
While LLM stories demonstrated promising potential as narrative interventions, some participants expressed concerns about the writing style and tone, which they felt diminished the overall impact. This issue arose from the use of certain phrasing or the participants’ familiarity with the way ChatGPT responds, making the stories feel less authentic. The tone of the responses was frequently mentioned, with some participants finding the tone to be too mechanical lacking the personal touch necessary for conversations around managing negative thoughts (Cheng et al., 2022). Others noted that the overly optimistic tone in certain stories did not always align with or acknowledge the severity of their struggles, which, in turn, made them feel that their challenges were being minimized (Haugeland et al., 2022).
@@ 5.2.3. Structured Reflection Activities
Participants in our study expressed appreciation for the reflection activity embedded within the stories, noting that such exercises helped them reevaluate their thoughts and emotions and potentially identify solutions to their problems. However, the reflection activity in our study was limited to a single question. Future studies might consider drawing inspiration from guided reflection activities, which involve a sequence of targeted questions focusing on specific aspects of one’s thoughts (O’Leary et al., 2018; Lee and Hong, 2017; Bhattacharjee et al., 2024b). These multi-question formats allow for a more structured presentation of thoughts, capturing nuances that a single question might miss. Incorporating such detailed questioning can enhance the benefits of LLM stories by providing deeper insights into their thoughts and challenges, and potentially helping find ways of solving them. (Bhattacharjee et al., 2024a; Sharma et al., 2023b; Park et al., 2021). However, it is important to note that answering these more detailed questions requires more time and effort from users. Therefore, careful consideration should be given to the timing and frequency of these questions to maintain a balance between depth of engagement and participant burden.
@@ 5.2.4. Improving Peer Support and Crowd-Powered Platforms
In the past, peer support and crowd-powered platforms have shown promising potential in providing wellbeing support, yet they face several inherent challenges that could be alleviated by an approach like ours. For example, the act of sharing and writing content can be daunting; LLMs can assist crowdworkers and peers by providing example stories and offering writing support to edit and restructure users’ own experiences (Kim et al., 2024; Reza et al., 2023; Gatos et al., 2021). Additionally, some individuals may be reluctant to open up to real people (Park et al., 2021). In these cases, the ability of LLMs to craft highly personalized narratives in a completely private manner — without the need to share personal details — could significantly increase their appeal, particularly for more inhibited users. Nonetheless, it is crucial to position LLMs as supportive tools that enhance the human-driven process and keep the essential human element in peer support intact.
@@ 5.2.5. Considerations for People’s Attitude Toward AI
While our study did not specifically measure people’s attitudes toward AI, participants’ comments suggest that such attitudes may influence their perceptions of LLM-enhanced stories. Many participants appeared open to the idea of AI-based tools providing support for managing negative thoughts, while others expressed resistance, reflecting a wide range of expectations regarding AI’s role in personal and emotional contexts (Kaya et al., 2024; Liu et al., 2022). This variation highlights the importance of considering how perceptions may shift depending on whether users are aware that a story is AI-generated or human-written — particularly since, in real-world deployments, authorship may not be masked.
As such, future research could benefit from directly measuring users’ attitudes toward AI (Grassini, 2023) and systematically investigating how knowledge of a story’s origin influences engagement, perceived helpfulness, and impact on negative thoughts and emotions. In addition, future DMH tools could allow users to express their attitudes toward AI during onboarding in order to adapt to those views. Gradually building rapport, rather than offering support immediately, may also help accommodate diverse expectations and comfort levels (O’Dell et al., 2022).
5.2.6. Considerations for Longitudinal Deployments
While our study explored user perceptions of narrative interventions as an SSI, deploying similar interventions across multiple sessions will introduce additional design considerations. Future longitudinal studies are essential to investigate these implications in greater depth, but our observations hint at several aspects that could be important to consider.
Repeated engagement with narratives, even when the content differs, may result in habituation (Rankin et al., 2009) if the stories consistently follow similar styles or structures, thereby reducing their effectiveness. To maintain user engagement, LLM-enhanced narratives should dynamically vary in format, length, tone, and writing style across sessions. Achieving this diversity may require insights from the literature on just-in-time adaptive interventions (Bhattacharjee et al., 2023b; Kabir et al., 2022), which emphasizes the role of contextual factors such as time, mood, emotions, and social context in shaping user engagement with DMH content. AI-driven systems could leverage this literature to create adaptive guidelines for LLMs, enabling the generation of contextually tailored stories. For instance, users might prefer shorter narratives during a busy workday but may engage more deeply with longer, reflective stories during a relaxed weekend (Bhattacharjee et al., 2023a). By integrating such dynamic adjustments, longitudinal interventions may be able to maintain user engagement and provide meaningful experiences that adapt to their changing needs and circumstances.
5.3. Limitations and Future Work
@@
Our study has several limitations that merit attention. First, our study population was confined to young adults. While some findings from this demographic might be generalizable to other groups, conducting similar studies with different populations could reveal new insights into their preferences and responses. Future studies should also consider including individuals with clinically elevated levels of depression and anxiety symptoms, keeping in mind the heightened need for risk management processes and oversight in such research.
Second, we did not embed specific action strategies within the narratives; rather, our goal was to communicate psychological principles for managing negative thoughts. Some participants mentioned that the stories inspired them to adopt a different mindset or consider taking small steps forward, but we did not systematically track whether or how participants acted on the stories. Future research should explore what kinds of actions, if any, individuals are motivated to take after engaging with such narratives. While our intervention did not include explicit suggestions, we recognize that introducing more action-oriented content in future systems may require careful consideration of timing and user readiness, as individuals may vary in their willingness or capacity to engage with such content depending on their context (Bhattacharjee et al., 2023b).
@@
Finally, our study relied primarily on user perceptions of the LLM-enhanced stories — their views on various qualities of the stories and their effectiveness as interventions. That is, we did not evaluate the objective qualities of the content itself. While we constrained the LLMs to work within the framework of previously established narratives supporting evidence-based psychological strategies, future research should assess adherence to these strategies and other inappropriate content that may appear in generated stories, potentially involving clinicians in the evaluation process to ensure the accuracy and appropriateness of the content.
@@ 6. Conclusion
LLMs offer promising capabilities to dynamically adapt and generate stories that can resonate with the unique challenges individuals face. Motivated by this potential, our study leveraged LLM to enhance and personalize existing human-written stories from past DMH interventions, comparing the LLM-enhanced versions with their original counterparts. We conducted an experiment with young adults, including crowdworkers and students, and observed that LLM stories were perceived to do better than human-written ones in conveying key takeaways, promoting reflection, and reducing belief in negative thoughts. These narratives were not only perceived as more relatable but also maintained a level of authenticity comparable to human-written stories. Our findings illuminate potential areas for future work on LLM-enhanced narrative interventions, such as balancing relatability with avoiding implausibility and refining the AI’s tone and writing style. We believe our work marks a significant step towards enhancing DMH interventions with emerging AI technology, enabling them to dynamically adapt based on user challenges.
@@ Acknowledgements.
We express our gratitude to Mary Czerwinski (Microsoft Research), Javier Hernandez (Microsoft Research), Haochen Song (University of Toronto), and Benjamin Kaveladze (Northwestern University) for their suggestions on relevant literature and their feedback on this work. This work was supported by grants from the National Institute of Mental Health (K01MH125172, R34MH124960), the Office of Naval Research (N00014-18-1-2755, N00014-21-1-2576), the Natural Sciences and Engineering Research Council of Canada (RGPIN-2019-06968), and the National Science Foundation (2209819). In addition, we acknowledge a gift from the Microsoft AI for Accessibility program to the Center for Behavioral Intervention Technologies that, in part, supported this work (http://aka.ms/ai4a). Ananya Bhattacharjee would also like to acknowledge support from the Inlight Research Fellowship, the Schwartz Reisman Institute Graduate Fellowship, and the Wolfond Scholarship in Wireless Information Technology.
Appendix A Non-LLM Stories and Corresponding LLM-Enhanced Examples
@@
Please answer this reflective question: Can you think of a time when reaching out to someone helped you break out of a rut?
@@
Relying on others may sometimes make you feel vulnerable, but sometimes it is in this vulnerability that we find our greatest strengths. There’s no shame in seeking help, after all, it’s okay to not have all the pieces together. We’re all a work in progress.
@@
I was thrust into independence at an early age, having realized that relying on others often led to disappointment. This developed a hardened shell around me, especially in romantic relationships where the fear of vulnerability reigned. I yearned for connection, yet the thought of opening up was paralyzing. Once I did open up, the outcome was a heartbreak that took me seven years to recover from. During those years, a shroud of chronic depression enveloped me. I pushed everyone away, building an even thicker protective barrier around myself.
@@
I feel too shy or anxious when I’m in social situations I feel overwhelmed and awkward in social situations, like I don’t know how to fit in or what to say. The anxiety just takes over, and I end up feeling invisible or out of place.
It’s exhausting, and I wish I could relax and be myself around others.
I am a shy person. It’s part of who I am. I sometimes feel bad about this, like I have lost out on experiences because of being shy and not showing my true self to the world. While I know I am as worthy of having a voice as anyone else, there’s this worry in my head that others will judge me.
@@
I’ve often found myself tangled in the webs of social anxiety, like a silent character adrift in a sea of voices. There’s this persistent whisper in my psyche, convincing me that everyone in the room is scrutinizing each slip, each stumble, and each stutter.
@@
Appendix B Narrative Qualities and Intervention Outcomes Across All Challenging Thoughts
In this section, we present participants’ evaluations of stories from both the LLM and Non-LLM conditions, categorized by which of the six challenging thoughts they addressed. Figures 2 through 5 illustrate their qualities as narratives (measures related to RQ1), and Figures 6 through 11 focus on intervention outcomes (measures related to RQ2). Each figure displays results separately for the crowdworker and student groups, with error bars representing the standard error of the mean and sample sizes indicated above each bar. In each of the figures, the six challenging thoughts correspond to the following:
  • Thought 1: I always feel really down on myself.
  • Thought 2: A lot of things could go wrong in the future.
  • Thought 3: I am struggling more than I should.
  • Thought 4: I haven’t achieved as much as other people have.
  • Thought 5: The people I have counted on have let me down.
  • Thought 6: I feel too shy or anxious when I’m in social situations.
Figure 2. Mean perceived authenticity across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived authenticity across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 3. Mean perceived ability to communicate key takeaways across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived ability to communicate key takeaways across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 4. Mean perceived ability to promote reflection across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived ability to promote reflection across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 5. Mean perceived balance of positivity with realistic struggles across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived balance of positivity with realistic struggles across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 6. Mean perceived reduction in belief in negative thoughts across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived reduction in belief in negative thoughts across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 7. Mean perceived reduction in emotion intensity across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived reduction in emotion intensity across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 8. Mean perceived relatability across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived relatability across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 9. Mean perceived capacity of inspiring solutions across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean perceived capacity of inspiring solutions across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 10. Mean likelihood of future engagement across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean likelihood of future engagement across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
Figure 11. Mean number of words written in response to the reflection question across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.
[…]
Mean number of words written in response to the reflection question across Non-LLM and LLM stories addressing six challenging thoughts, shown separately for crowdworkers (left) and students (right). Error bars represent the standard error of the mean, with sample sizes indicated above each bar.